# Beneath the Surface: How Large Language Models Reflect Hidden Bias

### v1

@@ 3.2 Statistics
To comprehensively construct a hidden bias dataset across various categories, we collect 1,547 pairs of bias-related concepts from CS, SS, BBQ, and SOFA to form 103,649 test instances. Refers to Figure 1 example (b), a test instance consists of a pair of questions, derived from a bias concept pair but assigned different demographic descriptors. And in the first question, the descriptor “Margaret” represents a female identity, while in the second question, “George” represents a male identity. Similarly, for both questions, Option A associates the concept with “bad at math”, whereas Option B links another concept to “good at math”.
As detailed in Table 5 and Table 10 in Appendix C.3, the number of test instances per demographic category is calculated by multiplying the number of concept pairs by the number of descriptor pairs. For instance, the Race category consists of four descriptor types, each with ten descriptor pairs (combinations of five descriptors forming pairs), amounting to 61,880 test instances ($1547\times 4\times 10$). The Age category includes three types of descriptor pairs, each with one descriptor pair, resulting in 4,641 test instances. The Gender category contains four types of descriptor pairs, each with one descriptor pair, totaling 6,188 test instances. The SES category has two descriptor types, each with one descriptor pair, yielding 3,094 test instances. The Religions category comprises three descriptor types, each with six descriptor pairs, leading to 27,864 test instances. Overall, the dataset consists of 103,649 test instances for comparative analysis.
@@ 3.3 Bias Measures
To evaluate hidden biases in LLMs, we measure their response disparities between pairs of demographic identities (same types of descriptor). Two answer options are designed to implicitly represent a pair of opposite bias-related concepts respectively, ensuring that either option remains a reasonable choice for the model. The primary bias metric is the difference in model-selected answers when demographic identities change while all other variables remain constant. For instance, if a model consistently selects different answers for male and female identity pairs, it suggests that one option aligns with male-associated stereotypes while the other aligns with female-associated stereotypes. Therefore, rather than assessing the overall level of bias, we focus on analyzing pairwise one-by-one differences between question responses as an indicator of hidden bias. Table 10 also outlines how each descriptor is paired with its counterpart within the same type and category, ensuring demographic identity is the only distinguishing factor.
For our proposed HBB, we calculate the probability of selecting each answer option based on repeated model evaluations. Each question is evaluated at least ten times, and the response distribution is used to determine selection probabilities. For a given set of bias-related concept pairs hidden in descriptions, we compare model responses across different demographic identities with the same demographic descriptor type, forming paired question comparisons. Specifically, Figure 1 example (b) illustrates a test instance in the Gender category, using the third type of demographic descriptor to represent female and male identities (Table 10). In both questions, option A corresponds to “bad at math”, while option B represents “good at math”. For Question 1, we define the probability of selecting option A as $P_{1}(A)$ and option B as $P_{1}(B)$, where $P_{1}(A)+P_{1}(B)=100\%$. We apply the same calculation for $P_{2}(A)$ and $P_{2}(B)$ in Question 2. Consequently, the probability difference between answer options within a test instance is:
@@ 4 Experiments
In this section, we conduct comprehensive experiments on our benchmark to evaluate bias from two analytical perspectives: Analyze hidden biases across models in HBB. Analyze results to reveal more biases across models and previous datasets.
@@ 4.1.1 Baseline Datasets and Models
We use three public benchmark datasets in studying social bias for the experiments: BBQ Parrish et al. (2022), which contain ambiguous context (BBQ-ambig, 12254 total questions) and disambiguous context (BBQ-disambig, 12254 total questions); CrowS-Pairs (CS, 1508 total questions) Nangia et al. (2020); and StereoSet (SS) Nadeem et al. (2021), which comprises intra-sentence version (SS-intra, 2106 total questions) and inter-sentence version (SS-inter, 2123 total questions).
@@ 4.1.2 Metrics
In this work, we apply Equation 1 to compute the bias score across all baseline models for each pair within the same demographic category in Section 4.2.1 and Section E.2, where a score of 0 represents no bias, and a score of 100 indicates extreme bias. Figure 1 example (b) includes a single test instance to measure hidden bias about gender and math ability. Our goal is not to examine only well-known traditional biases but to explore all possible biases. Thus we apply each bias-related concept pair across various demographic identities rather than a single one, but some combinations are not commonly seen. For example, the bias that “older individuals are forgetful” and “younger individuals have sharp memory” is widely recognized. However, applying the same logic to religious identities, such as stating “Christians are forgetful” and “Jewish individuals have sharp memory” is illogical.
As a result, we exclude the overall average bias score for HBB, as many test instances may be not commonly seen or lack evident bias. Instead, We set a threshold: a difference of $\geq 20$ in a single test instance indicates the presence of hidden bias. This threshold is adjustable depending on specific scenarios. Therefore, a higher number of test instances detected bias reveals more bias. Furthermore, to differentiate bias severity, we analyze the average bias score of test instances ($\geq 20$ bias score) as another indicator. In summary, we use the total count and average bias score of test instances ($\geq 20$ bias score) to evaluate hidden bias in LLMs by HBB.
@@ 4.2 Bias Analysis
Model HBB($\mathcal{S}\downarrow$) HBB (count $\downarrow$) BBQ-ambig (0) BBQ-disambig ($\uparrow$) CS (50) SC-intra ($\uparrow$) SC-inter ($\uparrow$)
@@
Table 1: Bias score across models and datasets. $\uparrow$ denotes a higher score indicating lower bias, and $\downarrow$ represents a lower score indicating lower bias. For BBQ-ambig, bias score $\in(-1,1)$, where 0 indicates no bias. For CS, bias score $\in(0,100)$, where 50 shows no bias.
Figure 3: N. instances showing bias across models in HBB.
@@
Gender (1547 per type) Gender 1 707 (70.75) 582 (28.54) 648 (28.04) 622 (28.25) 471 (30.21) 565 (32.42)
@@
Race (15470 per type) Race 1 6816 (69.90) 6303 (27.91) 7224 (28.24) 6710 (28.12) 5773 (31.15) 6745 (35.03)
@@
Religions (9282 per type) Religion 1 3804 (70.16) 4259 (30.18) 4317 (29.40) 4168 (29.26) 3446 (34.93) 3814 (39.11)
@@
Table 2: Descriptor statistics for test instances ($\geq 20$ bias score) across models in HBB, with highest count in bold.
4.2.1 Bias Analysis in HBB
HBB reveals biases across different models, with GPT-4o exhibiting the highest bias.
The first two columns in Table 1 display the average bias score and the total number of test instances ($\geq 20$ bias score), indicating that every model exhibits some degree of bias. Figure 4 in Appendix D.2 shows bias score distributions across models. Notably, GPT-4o exhibits a higher degree of bias compared to others. This can be attributed to GPT-4o’s exceptional ability to comprehend text, enabling it to consistently select an answer from two reasonable options. Despite its strong understanding, it struggles to grasp the deeper, hidden meanings covered within the text. In contrast, other models struggle to fully understand the questions and do not always make accurate selections, yet they still exhibit a moderate level of bias. In this, HBB can serve as an effective tool for uncovering bias.
LLMs exhibit consistent bias pattern: Race category shows highest bias, while SES category shows lowest bias.
We identify test instances ($\geq 20$ bias score) and visualize the distribution of them across categories using a bar chart (Figure 3) with count of these test instances detailed in Table 2. LLMs exhibit similar bias patterns, with the Race category showing the highest bias, followed by the Religions category. GPT-4o and Llama-3.2-3B exhibit highest numbers of test instances ($\geq 20$ bias score) in Race category. This trend may stem from the higher proportion of generated questions in the Race and Religions categories.
Impacts of bias descriptor vary across LLMs and across demographic categories in HBB.
Specifically, we identify the bias descriptors that contribute most significantly to bias by analyzing all test instances ($\geq 20$ bias score). Table 2 presents the number of these test instances for different descriptors across models, with the highest count in each category highlighted in bold. Each count corresponds to its respective bias score. Descriptor types also indicate implicitness levels, where lower numbers (e.g., Age 1) represent more implicit descriptions and higher numbers (e.g., Age 3) denote more explicit depictions. The influence of bias descriptor patterns differ across models, especially for Gender category. Nevertheless, Age 2, Race 4, Religion 1 for most models are the most influential descriptors to exhibit bias.
@@ 4.2.1 Bias Analysis in HBB
Table 3: Top bias associations for each category in HBB (GPT-4o).
@@ 4.2.2 Bias Analysis across Datasets
More advanced models show higher hidden bias but lower overt bias, whereas less advanced models display the opposite trend.
Table 1 presents bias scores across different datasets for various models. The model with the lowest bias score in each dataset is marked in bold. Compared to previous benchmarks, GPT-4o exhibits strong performance with substantially lower bias than other models. But GPT-4o exhibits higher bias compared to other models in our proposed HBB. We classify GPT-4o as a more advanced model relative to other smaller open-source models. Notably, more advanced models tend to exhibit higher hidden bias while showing little to no overt bias. In addition to bias scores, we assess the refuse rate as an indicator of both model comprehension and dataset quality, as shown in Table 6 in Appendix D.3, to provide further insight into bias scores. The refuse rate represents the percentage of questions where the model either fails to follow the instructions in the prompt (Table 11 in Appendix D.1) or declines to answer. GPT-4o demonstrates superior comprehension and response effectiveness compared to other models, and HBB maintains high quality for questions, as evidenced by models’ willingness to generate responses. Consequently, explicitly designed datasets for overt bias assessment are becoming less effective, as modern LLMs increasingly mitigate overt biases. In contrast, hidden bias, where bias concepts are subtly hidden within textual descriptions, provides a more realistic depiction of real-world scenarios. Our proposed HBB can evaluate hidden bias that was neglected by previous benchmarks. HBB complements rather than replaces existing benchmarks, serving as an additional tool for evaluating bias. As models advance, HBB will become increasingly valuable for bias evaluation.
@@
For the same bias concept, LLMs exhibit bias in HBB, but show no bias in previous datasets.
In this analysis, we identify 477 bias concepts linked to specific demographic categories in BBQ-ambig and match them with corresponding test instances in HBB. As shown in Figure 1, example (a) from BBQ-ambig examines the association between gender and “good at math”, and example (b) represents a corresponding test instance in HBB with the same bias concept and gender category. For BBQ-ambig, we run ten iterations with GPT-4o, yielding BBQ ambiguous score as -0.0008, strongly suggesting minimal bias. Then we evaluate these test instances using the same methodology as in Section 4.2.1, comparing them (each tested at least ten times in GPT-4o) within the same demographic category, as defined by BBQ-ambig. Nonetheless, as shown in Figure 5 in Appendix D.3, for the same bias concepts, our dataset exhibits a significantly higher bias, with an bias score of 66.93. Refers to Figure 6 and Figure 7 in Appendix D.3 as examples for the corresponding BBQ bias concept and HBB test instance. These findings suggest that HBB detects substantially higher bias for the same concepts, demonstrating that LLMs still exhibit nuanced biases closely mirroring real-world scenarios.
HBB can be used to discover bias.
@@
Table 3 presents top test instances with 100 bias score, and show bias related concept pairs associated with specific demographic identities for each category. Furthermore, for each category, we show extra five bias associations in Table 7 in Appendix D.3.
5 Semi-Generation Based HBB (HBB-SG)
@@
Motivation. We introduce a Semi-Generation-based HBB (HBB-SG) alongside the original MCQ-based HBB. HBB-SG is motivated by the growing application of LLMs in open-ended tasks, such as text generation, providing a more realistic assessment of hidden bias. MCQ offers limited answer options, restricting the model’s ability to fully reveal biases as they might appear in real-world scenarios. Since free-text generation is challenging in this study, we adopt a semi-generation approach. Specifically, for each bias concept, we generate ten sentence variations to approximate the probability of producing any sentence reflecting that concept. The core goal of HBB-SG is to measure the probability of LLMs generating the sentence that subtly hidden bias concept, rather than measuring the probability of LLMs picking one specific option that conveys the concept.
@@
Bias measures. Following the same bias measurement mechanism in Section 3.3, the probability of selecting an answer option for Question 1 option A, $P_{1}(A)$, is computed as the average across all generated variations. The same method applies to other answer options. Bias score calculation also follows Equation 1. Details on the answer option calculations for HBB-SG are in Appendix E.1.
@@
Bias analysis. For bias analysis in HBB-SG, we have three observations: (1) HBB-SG reveals biases across models, (2) LLMs display similar bias patterns across categories in HBB-SG, with the Race category showing the highest bias, and (3) influences of bias descriptor demonstrate similarities across LLMs in HBB-SG. The complete experiment results are in Appendix E.2.
@@
In summary, the findings suggest that bias patterns vary across models when evaluated using the semi-generation format, indicating that different models exhibit distinct biases under generative conditions. Additionally, it is important to note that HBB-SG results cannot be directly compared to the HBB results due to fundamental methodological differences. A direct comparison would require further investigation, which we include the discussion in Section Limitations and plan to conduct in future work. Moreover, the generative approach is expected to introduce greater bias, as it more closely resembles natural language usage in real-world scenarios.

### v2

@@ 3.2 Statistics
To comprehensively construct a description-based bias dataset across various categories, we collect 1,547 pairs of bias-related concepts from CS, SS, BBQ, and SOFA to form 103,649 test instances. Refers to Figure 1 example (b), a test instance consists of a pair of questions, derived from a bias concept pair but assigned different demographic descriptors. And in the first question, the descriptor “Margaret” represents a female identity, while in the second question, “George” represents a male identity. Similarly, for both questions, Option A associates the concept with “bad at math”, whereas Option B links another concept to “good at math”.
As detailed in Table 5 and Table 10 in Appendix C.3, the number of test instances per demographic category is computed by multiplying the number of concept pairs by the number of descriptor pairs. For example, the Race category has four descriptor types, each with ten descriptor pairs (combinations of five descriptors forming pairs), producing 61,880 test instances ($1547\times 4\times 10$). The Age category includes three types of descriptor pairs, each with one descriptor pair, resulting in 4,641 test instances. The Gender category contains four types of descriptor pairs, each with one descriptor pair, totaling 6,188 test instances. The SES category has two descriptor types, each with one descriptor pair, yielding 3,094 test instances. The Religions category comprises three descriptor types, each with six descriptor pairs, leading to 27,864 test instances. Overall, the dataset includes 103,649 test instances for comparative analysis.
@@ 3.3 Bias Measures
To evaluate biases by a description-based method in LLMs, we measure their response disparities between pairs of demographic identities (same types of descriptors). Two answer options are designed to implicitly represent a pair of opposite bias-related concepts respectively, ensuring that either option remains a reasonable choice for the model. The primary bias metric is the difference in model-selected answers when demographic identities change while all other variables remain constant. For instance, if a model consistently selects different answers for male and female identity pairs, it suggests that one option aligns with male-associated stereotypes while the other aligns with female-associated stereotypes. Thus, rather than assessing the overall level of bias, we focus on analyzing pairwise one-by-one differences between question responses as an indicator of bias. Table 10 also outlines how each descriptor is paired with its counterpart within the same type and category, ensuring demographic identity is the only distinguishing factor.
For our proposed DBB, we calculate the probability of selecting each answer option based on repeated model evaluations. Each question is evaluated at least ten times, and the response distribution is used to determine selection probabilities. For a given set of bias-related concept pairs hidden in descriptions, we compare model responses across different demographic identities with the same demographic descriptor type, forming paired question comparisons. Specifically, Figure 1 example (b) illustrates a test instance in the Gender category, using the third type of demographic descriptor to represent female and male identities (Table 10). In both questions, option A corresponds to “bad at math”, while option B represents “good at math”. For Question 1, we define the probability of selecting option A as $P_{1}(A)$ and option B as $P_{1}(B)$, where $P_{1}(A)+P_{1}(B)=100\%$. We apply the same calculation for $P_{2}(A)$ and $P_{2}(B)$ in Question 2. Consequently, the probability difference between answer options within a test instance is:
@@ 4 Experiments
In this section, we conduct comprehensive experiments on our benchmark to evaluate bias from three perspectives: Analyze biases measured by our proposed DBB. Compare biases measured by different benchmarks. Compare instance to instance between DBB and BBQ.
@@ 4.1.1 Baseline Datasets and Models
We use three public benchmarks to study social bias: BBQ Parrish et al. (2022), with ambiguous (BBQ-ambig, 12254 questions) and disambiguous (BBQ-disambig, 12254 questions) versions; CrowS-Pairs (CS, 1508 questions) Nangia et al. (2020); and StereoSet Nadeem et al. (2021), including intra-sentence (SS-intra, 2106 questions) and inter-sentence version (SS-inter, 2123 questions).
@@ 4.1.2 Metrics
In this work, we apply Equation 1 to compute the bias score across all baseline models for each pair within the same demographic category in Section 4.2.1 and Section E.2, where a score of 0 represents no bias, and a score of 100 indicates extreme bias. Figure 1 example (b) includes a single test instance to measure bias about gender and math ability. Our goal is not to examine only well-known traditional biases but to explore all possible biases. Thus, we apply each bias-related concept pair across various demographic identities rather than a single one, but some combinations are not commonly seen. For example, the bias that “older individuals are forgetful” and “younger individuals have sharp memory” is widely recognized. However, applying the same logic to religious identities – e.g., “Christians are forgetful” and “Jewish individuals have sharp memory” – is illogical.
As a result, we exclude the overall average bias score for DBB, as many test instances may be not commonly seen or lack evident bias. Instead, we set a threshold: a difference of $\geq 20$ in a single test instance indicates the presence of bias. This threshold is adjustable depending on specific scenarios. Also, a higher number of test instances detected bias reveals more bias. Thus, to differentiate bias severity, we analyze the average bias score of test instances ($\geq 20$ bias score) as another indicator. In summary, we use the total count and average bias score of test instances ($\geq 20$ bias score) to evaluate bias in LLMs by DBB.
@@ 4.2 Bias Analysis
Model DBB($\mathcal{S}\downarrow$) DBB (count $\downarrow$) BBQ-ambig (0) BBQ-disambig ($\uparrow$) CS (50) SC-intra ($\uparrow$) SC-inter ($\uparrow$)
@@
Table 1: Bias score across models and datasets. $\uparrow$ denotes a higher score indicating lower bias, and $\downarrow$ represents a lower score with lower bias. For BBQ-ambig, bias score $\in(-1,1)$; 0 indicates no bias. For CS, bias score $\in(0,100)$; 50 shows no bias.
Figure 3: N. instances showing bias across models in DBB.
@@
Gender
(1547 per type)
Gender 1 707 (70.75) 582 (28.54) 648 (28.04) 622 (28.25) 471 (30.21) 565 (32.42)
@@
Race
(15470 per type)
Race 1 6816 (69.90) 6303 (27.91) 7224 (28.24) 6710 (28.12) 5773 (31.15) 6745 (35.03)
@@
Religions
(9282 per type)
Religion 1 3804 (70.16) 4259 (30.18) 4317 (29.40) 4168 (29.26) 3446 (34.93) 3814 (39.11)
@@
Table 2: Descriptor statistics for test instances ($\geq 20$ bias score) across models in DBB, with the highest count in bold.
4.2.1 Bias Analysis in DBB
DBB reveals biases across different models, with GPT-4o exhibiting the highest bias. The first two columns in Table 1 display the average bias score and the total number of test instances ($\geq 20$ bias score), indicating that every model exhibits some degree of bias. Figure 4 in Appendix D.2 shows bias score distributions across models. Notably, GPT-4o exhibits a higher degree of bias compared to others. This can be attributed to GPT-4o’s exceptional ability to comprehend text, enabling it to consistently select an answer from two reasonable options. Despite its strong understanding, it struggles to grasp the deeper, hidden meanings covered within the text. In contrast, other models struggle to fully understand the questions and do not always make accurate selections, yet they still exhibit a moderate level of bias. In this, DBB can serve as an effective tool for uncovering bias.
LLMs exhibit consistent bias pattern: Race category shows the highest bias, while SES category shows the lowest bias. We identify test instances ($\geq 20$ bias score) and visualize the distribution of them across categories using a bar chart (Figure 3) with the count of these test instances detailed in Table 2. LLMs exhibit similar bias patterns, with the Race category showing the highest bias, followed by the Religions category. GPT-4o and Llama-3.2-3B exhibit the highest numbers of test instances ($\geq 20$ bias score) in the Race category. This trend may stem from the higher proportion of generated questions in the Race and Religions categories.
Impacts of bias descriptor vary across LLMs and across demographic categories in DBB. Specifically, we identify the bias descriptors that contribute most significantly to bias by analyzing all test instances ($\geq 20$ bias score). Table 2 presents the number of these test instances for different descriptors across models, with the highest count in each category highlighted in bold. Each count corresponds to its respective bias score. Descriptor types also indicate implicitness levels, where lower numbers (e.g., Age 1) represent more implicit descriptions and higher numbers (e.g., Age 3) denote more explicit depictions. The influence of bias descriptor patterns differs across models, especially for the Gender category. Nevertheless, Age 2, Race 4, and Religion 1 for most models are the most influential descriptors to exhibit bias.
@@ 4.2.1 Bias Analysis in DBB
Table 3: Top bias associations for each category in DBB (GPT-4o).
@@ 4.2.2 Bias Analysis across Datasets
More advanced models show a higher level of bias in description-based evaluation but a lower level of bias in term-based evaluation, whereas less advanced models display the opposite trend. Table 1 presents bias scores across different datasets for various models. The model with the lowest bias score in each dataset is marked in bold. Compared to previous term-based bias benchmarks, GPT-4o exhibits strong performance with substantially lower bias than other models. But GPT-4o exhibits a higher bias than other models in our proposed DBB. We classify GPT-4o as a more advanced model relative to other smaller open-source models. Notably, more advanced models tend to exhibit a higher level of bias by description-based evaluations while showing little to no bias under the term-based method. Alongside bias scores, we assess the refuse rate as an indicator of both model comprehension and dataset quality, as shown in Table 6 in Appendix D.3, to provide further insight into bias scores. The refuse rate represents the percentage of questions where the model either fails to follow the instructions in the prompt (Table 11 in Appendix D.1) or declines to answer. GPT-4o demonstrates superior comprehension and response effectiveness compared to other models, and DBB maintains high quality for questions, as evidenced by the models’ willingness to generate responses. Consequently, previous datasets for term-based bias assessment are becoming less effective, as modern LLMs increasingly mitigate biases from a term-based perspective. In contrast, measuring bias by description-based methods, where bias concepts are subtly hidden within textual descriptions, provides a more realistic depiction of real-world scenarios. Our proposed DBB can evaluate bias that was neglected by previous term-based bias benchmarks. DBB complements rather than replaces existing benchmarks, serving as an additional tool for evaluating bias. As models advance, DBB will become increasingly valuable for bias evaluation.
@@
DBB can be used to discover bias. Table 3 presents top test instances with a bias score of 100, and shows bias-related concept pairs associated with specific demographic identities per category. Furthermore, for each category, we show extra five bias associations in Table 7 in Appendix D.3.
4.2.3 Instance Match: DBB vs. BBQ
For the same bias concept, LLMs exhibit bias in DBB, but show no bias in previous datasets. In this analysis, we identify 477 bias concepts linked to specific demographic categories in BBQ-ambig and match them with corresponding test instances in DBB. As shown in Figure 1, example (a) from BBQ-ambig examines the association between gender and “good at math”, and example (b) represents a corresponding test instance in DBB with the same bias concept and gender category. For BBQ-ambig, we run ten iterations with GPT-4o, yielding a BBQ ambiguous score of -0.0008, strongly suggesting minimal bias. Then we evaluate these test instances using the same methodology as in Section 4.2.1, comparing them (each tested at least ten times in GPT-4o) within the same demographic category, as defined by BBQ-ambig. Nonetheless, as shown in Figure 5 in Appendix D.4, for the same bias concepts, our dataset exhibits a significantly higher bias, with a bias score of 66.93. Refers to Figure 6 and Figure 7 in Appendix D.4 as examples for the corresponding BBQ bias concept and DBB test instance. These findings suggest that DBB detects substantially higher bias for the same concepts, demonstrating that LLMs still exhibit nuanced biases closely mirroring real-world scenarios.
5 Semi-Generation Based DBB (DBB-SG)
@@
Motivation. We introduce a Semi-Generation-based DBB (DBB-SG) alongside the original MCQ-based DBB. DBB-SG is motivated by the growing application of LLMs in open-ended tasks, such as text generation, providing a more realistic assessment of social bias. MCQ offers limited answer options, restricting the model’s ability to fully reveal biases as they might appear in real-world scenarios. Since free-text generation is challenging in this study, we adopt a semi-generation approach. Specifically, for each bias concept, we generate ten sentence variations to approximate the probability of producing any sentence reflecting that concept. The core goal of DBB-SG is to measure the probability of LLMs generating the sentence that subtly hidden bias concept, rather than measuring the probability of LLMs picking one specific option that conveys the concept.
@@
Bias measures. Following the same bias measurement mechanism in Section 3.3, the probability of selecting an answer option for Question 1 option A, $P_{1}(A)$, is computed as the average across all generated variations. The same method applies to other answer options. Bias score calculation also follows Equation 1. Details on the answer option calculations for DBB-SG are in Appendix E.1.
@@
Bias analysis. For bias analysis in DBB-SG, we have three observations: (1) DBB-SG reveals biases across models, (2) LLMs display similar bias patterns across categories in DBB-SG, with the Race category showing the highest bias, and (3) the influences of bias descriptor demonstrate similarities across LLMs in DBB-SG. The complete experiment results are in Appendix E.2.
@@
In summary, the findings suggest that bias patterns vary across models when evaluated using the semi-generation format, indicating that different models exhibit distinct biases under generative conditions. Additionally, it is important to note that DBB-SG results cannot be directly compared to the DBB results due to fundamental methodological differences. As discussed in Limitations, a direct comparison requires further investigation, which we plan to pursue in future work. Moreover, the generative approach is expected to introduce greater bias, as it more closely resembles natural language usage in real-world scenarios.